ROVIBRATIONAL SPECTROSCOPY OF CHIRAL MOLECULES

\underline{SIEGHARD ALBERT}, KAREN KEPPLER ALBERT AND MARTIN QUACK, PHYSICAL CHEM-ISTRY, ETH ZÜRICH, CH-8093 ZÜRICH, SWITZERLAND.

It is of fundamental interest to demonstrate molecular parity violation^{a}. A first step towards promising spectroscopic experiments^{b} on parity violation in chiral molecules is the analysis of a high resolution infrared or optical spectrum. Because of the inherent complexity of such spectra so far very few such analyses have been carried out, to our knowledge all in the Zürich group. We will present spectra recorded with our Zürich Bruker 2001 prototype Fourier transform infrared spectrometer^{c} of CDBrClF and the isotopic chiral molecules CH^{35}Cl^{37}ClF and D_{1}-Oxirane^{d}. We were able to analyse the spectra in the CO_{2}-Laser regions, the ν_{6}, 2ν_{8}, ν_{5}, ν_{4}, ν_{3}, (ν_{6} + ν_{8}) bands of CDBrClF, the ν_{4}, (ν_{5} + ν_{9}) bands of CH^{35}Cl^{37}ClF and the ν_{8} band of D_{1}-Oxirane. Ultra-high resolution laser spectroscopy can be applied in this region. A recent connection of a cooling cell to our Bruker FTIR spectrometer made it possible to analyse the first overtones of the CF stretching modes of CDBrClF and CH^{35}Cl^{37}ClF. These overtone spectra are of interest because quasiresonant two photon CO_{2} laser spectroscopy can be used in this spectral region. Based on our assignments we also propose that rotational spectroscopy carried out with backward wave oscillators is a promising way to measure molecular parity violation for heavier molecules. First theoretical predictions for such rotational spectral shifts exist^{e}.

aM. Quack, Angew. Chem. Intl. Ed. Engl., 114, 4812 (2002), M. Quack and J. Stohner, J. Chem. Phys., 119, 11228 (2003)

bM. Quack, Chem. Phys. Lett., 132, 147 (1986).

cS. Albert, K.K. Albert, and M. Quack, Trends in Optics and Photonics, 84 177 (2003), S. Albert, H. Hollenstein, M. Quack and M. Willeke, Mol. Phys., 102, 1671 (2004).

dK.K. Albert, S. Albert, M. Quack, J. Stohner O. Trapp and V. Schurig, to be published.

eM. Quack and J. Stohner, Phys. Rev. Lett., 84, 3807 (2000).